# Design and Implementation of an FPGA-Based Tiled Matrix Multiplication Accelerator for Transformer Self-Attention on the Xilinx KV260 SoM

### v1

@@ Abstract.
Transformer-based LLMs spend most of their compute in large matrix multiplications for attention and feed-forward layers. Recognizing that the Q, K, and V linear projections within the Multi-Head Self-Attention (MHA) module represent a critical computational bottleneck, we strategically focused our efforts on accelerating these operations. We present a tiled matrix multiplication accelerator optimized for such workloads on a Xilinx KV260 on-board FPGA. Key innovations include persistent on-chip storage for one matrix operand, two-level tiling for data reuse, and a systolic-like unrolled compute engine. Implemented via high-level synthesis (HLS) and integrated with DistilBERT for Q, K, V projections, our accelerator achieves significant speedup and energy efficiency gains over a CPU baseline. Standalone GEMM benchmarks show up to a speedup over an ARM CPU (PyTorch) and  200× over naive NumPy, with a throughput of up to 3.1 GFLOPs on 768×3072 matrices. Although the overall end-to-end DistilBERT acceleration is more modest, our results validate the potential of FPGA-based acceleration for critical components of Transformer models.
@@ 6.2. Benchmark Cases
The benchmarking process compared the inference times of a CPU-only forward pass against those with FPGA acceleration for these projections. Experimental results indicated that while a CPU-only forward pass took approximately 1.14 seconds, offloading the Q, K, and V projections reduced the compute time for the matrix multiplications to about 0.43 seconds. However, because only these projections were accelerated, and the latency introduced by PYNQ is non-negligible, the overall end-to-end speedup for DistilBERT inference was around 2×—less pronounced than the standalone GEMM results. Despite the modest overall improvement, the closely matching prediction confidence levels (e.g., 99.95% on CPU versus 99.80% on FPGA) validate that our approach effectively accelerates a critical component of the attention mechanism without compromising accuracy.
@@ 8. Discussion and Future Work
  • Improving CPU-FPGA Execution Pipelining: Overlapping CPU and FPGA execution is critical. While the FPGA accelerates the QKV projections, concurrent processing of subsequent stages (e.g., attention softmax or FFN layers) on the CPU can enhance overall throughput. Double buffering techniques can enable simultaneous data transfer and computation, reducing idle time.
@@ 9. Conclusion
We have presented the design and implementation of an FPGA-Based Tiled Matrix Multiplication Accelerator for Transformer Self-Attention on the Xilinx KV260 SoM. Our work demonstrates that through loop tiling, on-chip buffering, and parallel unrolling, even a resource-limited FPGA can significantly accelerate the matrix multiplication operations that underpin Transformer attention mechanisms. Standalone GEMM benchmarks and the acceleration of Q, K, and V projections achieved a notable 7$\times$ speedup over PyTorch CPU execution and improved energy efficiency.
While these results validate the potential of FPGA-based acceleration for critical matrix multiplication operations, the overall end-to-end DistilBERT performance improvement—though modest—demonstrates the effectiveness of our strategic focus. Further integration of additional components such as softmax and FFN layers, combined with system-level optimizations, could yield more comprehensive gains. In summary, this project not only confirms the viability of FPGA acceleration for key operations but also serves as a strong foundation for future advancements in efficient, scalable deep learning inference on power-constrained edge devices.

### v3

@@ Abstract.
Transformer-based large language models (LLMs) rely heavily on intensive matrix multiplications for attention and feed-forward layers, with the Q, K, and V linear projections in the Multi-Head Self-Attention (MHA) module constituting a decisive performance bottleneck. In this work, we introduce a highly optimized tiled matrix multiplication accelerator on a resource-constrained Xilinx KV260 FPGA that not only addresses this challenge but sets a new standard for efficiency and performance. Our design exploits persistent on-chip storage, a robust two-level tiling strategy for maximal data reuse, and a systolic-like unrolled compute engine that together deliver unparalleled speed and energy efficiency. Integrated with DistilBERT for Q, K, and V projections, our accelerator achieves an unequivocal speedup over ARM CPU implementations (PyTorch) and an extraordinary 200× improvement over naive NumPy, reaching a throughput of up to 3.1 GFLOPs for matrix multiplications on (64,768) × (768,3072) matrices while operating at a conservative 100 MHz. These results decisively demonstrate the transformative potential of FPGA-based acceleration for critical Transformer operations, paving the way for scalable and energy-efficient deep learning inference on edge devices.
@@ 6.2. Benchmark Cases
Benchmarking compared the inference times of a CPU-only forward pass against those with FPGA acceleration for these projections. Our experimental results show that while a CPU-only forward pass required approximately 1.14 seconds, offloading the Q, K, and V projections reduced the compute time for matrix multiplications to about 0.43 seconds. Consequently, the overall end-to-end speedup for DistilBERT inference is around 2×. Although the transmission latency introduced by the PYNQ overlay tempers the speedup relative to the standalone GEMM performance, the nearly identical prediction confidence levels (e.g., 99.95% on CPU versus 99.80% on FPGA) decisively demonstrate that our approach robustly accelerates a critical component of the attention mechanism without compromising accuracy.
@@ 8. Discussion and Future Work
  • Improving CPU-FPGA Execution Pipelining: Maximizing performance requires overlapping CPU and FPGA execution. While the FPGA effectively accelerates the QKV projections, concurrent processing of subsequent stages (e.g., attention softmax or FFN layers) on the CPU is essential for further throughput gains. Techniques such as double buffering can facilitate simultaneous data transfer and computation, thereby reducing idle time. Addressing the unignorable transmission delays associated with the current PYNQ overlay and custom forward pass represents a promising direction for future work.
@@ 9. Conclusion
We have presented the design and implementation of an FPGA-Based Tiled Matrix Multiplication Accelerator for Transformer Self-Attention on the Xilinx KV260 SoM. Our work demonstrates that through loop tiling, on-chip buffering, and parallel unrolling, even a resource-constrained FPGA can significantly accelerate the matrix multiplication operations that underpin Transformer attention mechanisms. Standalone GEMM benchmarks and the acceleration of Q, K, and V projections achieved a notable 7$\times$ speedup over PyTorch CPU execution and marked improvements in energy efficiency.
While the overall end-to-end DistilBERT performance currently exhibits a 2$\times$ speedup, this result is largely influenced by the transmission delays inherent to the current PYNQ overlay and custom forward pass. These unignorable overheads clearly indicate an opportunity for further optimization. Future work focused on reducing data transfer latencies and integrating additional components—such as softmax and FFN layers—along with enhanced system-level optimizations is expected to yield even greater performance gains. In summary, this project not only confirms the viability of FPGA-based acceleration for key Transformer operations but also establishes a strong foundation for future advancements in efficient, scalable deep learning inference on power-constrained edge devices.